OpenStudy (anonymous):

Help please! Lim ->1 Square Root of((x)- (x^2)/(1- sqare root of x)

OpenStudy (anonymous):

\[\frac{x-x^2}{1-\sqrt{x}}\]?

OpenStudy (anonymous):

\[\frac{x-x^2}{1-\sqrt{x}}=\frac{x(1-x)}{1-\sqrt{x}}=\frac{x(1-\sqrt{x})(1+\sqrt{x})}{1-\sqrt{x}}\] \[=x(1+\sqrt{x})\] then put in \(1\) where you see \(x\)

OpenStudy (anonymous):

Sorry its (sqrt of x) -x^2 for the numerator